OpenStudy (anonymous):

please help will medal use the given array to solve the problem and explain your thinking simplify fractions if needed 4over 5 x 25=? ***** ***** ***** ***** *****

OpenStudy (anonymous):

what is the array?

OpenStudy (anonymous):

Did you mean\[\frac{ 4 }{ 5 } * 25\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok well think of it like this \[\frac{ 4 }{ 5 } * \frac{ 25 }{ 1 }\] Then just multiply to get\[\frac{ 4*25 }{ 5*1}=\frac{ 100 }{ 5 }\] Then you just simplify to get \[\frac{ 20 }{ 1 }\] Which is equal to just 20. So \[\frac{ 4 }{ 5 } * \frac{ 25 }{ 1 }=20\]

OpenStudy (anonymous):

26 stars is the array

OpenStudy (anonymous):

25 not 26

OpenStudy (anonymous):

So what exactly are you looking for? Is 20 not the answer?

OpenStudy (anonymous):

what? do you mean?

OpenStudy (anonymous):

would 20 be the answer?

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

If you are looking for the answer of \[\frac{ 4 }{ 5 }*25\] then yes, 20 is the answer.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

4* 1 *12 _ 2

OpenStudy (anonymous):

So the question is \[4*\frac{ 1 }{ 2 }*12=?\] And you want to know what "?" is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well the first thing to do is just \[4*12\]which equals 48 So now you have \[\frac{ 1 }{ 2 }*48\]which can be re-written to fraction form to look like\[\frac{ 1 }{ 2 }*\frac{ 48 }{ 1 }\to \frac{ 1*48 }{ 2*1 }=\frac{ 48 }{ 2 }\] Then divide 48 by 2 to get\[\frac{ 24 }{ 1 }=24\] So your answer is 24

OpenStudy (anonymous):

i have another

OpenStudy (anonymous):

Ok go for it

OpenStudy (anonymous):

3 _ 9 x 12 =?

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

are you a girl legendarysadist?

OpenStudy (anonymous):

\[\frac{ 3 }{ 9 }*\frac{ 12 }{ 1 }\to \frac{ 3*12 }{ 9*1 }\to \frac{ 36 }{ 9 }=4\]So your answer is 4

OpenStudy (anonymous):

And no, I'm a guy

OpenStudy (anonymous):

oh you have a pic of a girl so yeah honest mistake

OpenStudy (anonymous):

im sorry

OpenStudy (anonymous):

lol it's fine

OpenStudy (anonymous):

ps i still need help with 2 more can you help?

OpenStudy (anonymous):

Sure, I can still help

OpenStudy (anonymous):

5x 3 =? _ 5

OpenStudy (anonymous):

\[\frac{ 5 }{ 1 }*\frac{ 3 }{ 5}\to \frac{ 5*3 }{ 1*5 }\] Do you think you can take it from this point?

OpenStudy (anonymous):

um idk what is that? i don't understand

OpenStudy (anonymous):

Well what is 5 times 3? And what is 1 times 5?

OpenStudy (anonymous):

15 and 1

OpenStudy (anonymous):

there look

OpenStudy (anonymous):

Close, it's 15 and 5. Anything times 1 is the other number. Do you know 15/5=?

OpenStudy (anonymous):

15x5?

OpenStudy (anonymous):

15 divided by 5. In other words, how many times will 5 go into 15?

OpenStudy (anonymous):

yes and its 3

OpenStudy (anonymous):

Good job! 3 is the answer to the problem.

OpenStudy (anonymous):

oh wow thank you i have just one more

OpenStudy (anonymous):

use the numbers 4,5 and 35 to write this problem with two different factors that result in the same product of 28